OpenStudy (anonymous):

A trajectory has an initial velocity of 1,000 meters per second, and 5 seconds later, it has a velocity of 500 meters per second. Which is the acceleration of the trajectory?

OpenStudy (shamim):

If u throw ur projectile vertically upward, then required formuls will b v=u-at

OpenStudy (shamim):

Given v=500m/s u=1000m/s t=5sec a=?